the sum of the digit of a two digit number is 11. when the digits are reversed, the new number is 24 less than four time the original number. what is the original number?

Answer by mananth(16949)

the sum of the digit of a two digit number is 11. when the digits are reversed, the new number is 24 less than four time the original number. what is the original number?
...
let the ten's digit be x
units digit be y
..
x+y =11........................1
10y+x=4(10x+y)-24
10y+x=40x+4y-24
6y-39x=-24.....................2
multiply (1) by -6
-6x-6y=-66
add it to(2)
-6x-39x-6y+6y=--66-24
-45x=-90
/-45
x=2
...
x+y=11
plug x
2+y=11
y=9
So the number is 29
...
m.ananth@hotmail.ca